Sound waves from a loudspeaker spread nearly uniformly in all directions if the wavelength of the sound is much larger than the diameter of the loudspeaker. (a) Calculate the frequency for which the wavelength of sound in air is ten times the diameter of the speaker if the wavelength is much shorter than the diameter of the speaker. Calculate the frequency a which the wavelength of the sound is one tenth of the diameter of the speaker described above. Take the speed of sound to be 340 m s-1.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. According to the question, a) λ = 20 cm × 10 = 200 cm = 2 m v = 340 m/s so, n = v/ λ = 340/2 = 170 Hz. N = v/λ ⇒ 340/2 * 10^2 = 17.000 Hz = 17 KH base 2 (because ⇒ = 2 cm = 2 * 10^2 m)